The mean of 4 numbers is 12.The mean of 3 is 14.What is the value of the 4th number
1
Expert's answer
2020-06-02T18:59:50-0400

The mean of 4 numbers is 12


"{a_1+a_2+a_3+a_4\over 4}=12"

"a_4=12(4)-(a_1+a_2+a_3)"

The mean of 3 numbers is 14


"{a_1+a_2+a_3\over 3}=14"

"a_1+\na_2+a_3=14(3)"

Then


"a_4=48-42=6"

 The value of the 4th number is "6."